OpenStudy (anonymous):

Solve the given equation. Check your solution. 3|2a + 7| =3a + 12

OpenStudy (anonymous):

6a+21=3a+12 3a=-9 a=-3

OpenStudy (anonymous):

divide by 3 first

OpenStudy (anonymous):

wait I'm not sure

OpenStudy (anonymous):

yeah, that is not right

OpenStudy (bahrom7893):

3(2a+7)=3a+12 -3(2a+7)=3a+12

OpenStudy (anonymous):

yeah I forgot to split it

OpenStudy (bahrom7893):

Lost connection -_-

OpenStudy (anonymous):

start with \[|2a+7|=a+4\]

OpenStudy (bahrom7893):

^lol yea, that's a pretty good advice

OpenStudy (anonymous):

then if \(a>-\frac{7}{2}\) you have \(|2a+7|=2a+7\) solve \[2a+7=a+4\] in two steps

OpenStudy (anonymous):

if \(a<-\frac{7}{2}\) you have \(|2a+7|=-2a-7\) solve \[-2a-7=a+4\] again in two steps